10. The density of water at 200C is 1000kgm-3 and its acoustic impedance is 1.5 x 106kgm-2s-1. What is the speed of sound ,c, in water at this temperature


1
Expert's answer
2021-03-26T11:40:46-0400

The acoustic impedance is


"z=\rho c,\\\space\\
c=\frac z\rho=\frac{1.5·10^6}{1000}=1500\text{ m/s}."